Based on a​ survey, assume that 44​% of consumers are comfortable having drones deliver their purchases. Suppose that we want to find the probability that when six consumers are randomly​ selected, exactly four of them are comfortable with delivery by drones. Identify the values of​ n, x,​ p, and q.
Answer #1

Let X denotes the number of consumers  comfortable with delivery by drones.

Then,

X will follow binomial distribution with parameters n=6 and p = 44% = 0.44

n =6

p = 0.44

x = 4

q = 1-p = 1- 0.44 = 0.56

The probability exactly four of them are comfortable with delivery by drones, P(X=4)

P(X = 4) = (6) +0.44* 0.566-4

(3) 0.4 0.56

= 0.1763